let z be a complex number such that |z|=5 .Find the maximum value of |z+3+4i|

Grade:11

1 Answers

Arun
6 years ago
We know from triangle inequality that|z+3+4i|‹=|z|+ |3+4i|Hence |z+3+4i|‹= 5 + 5|z+3+4i|‹= 10Hence maximum value of |z+3+4i| will be 10